OpenStudy (anonymous):

I WILL GIVE MEDAL AND BECOME FAN A community group is planning the expansion of a square flower garden in a city park. If each side of the original garden is increased by 9 meters, the new total area of the garden will be 169 square meters. Find the length of each side of the original garden. A. 9 m B. 13 m C. 2 m D. 4 m Please tell me the answer and how you got it Thanks!

OpenStudy (anonymous):

169 squared is 13 and the old garden was increased by 9 so 13-9=4m

OpenStudy (anonymous):

Ohhhhh I'm a dummy >.< Thanks so much for helping! Can you help me with a little bit more please? :)

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Solve the equation using square roots. 5x^2-30=0 A. x=6,-6 B. x=Square Root of 6, Square Root of -6 C. x=3, -3 D. No solutions

OpenStudy (anonymous):

Simplifying 5x2 + -30 = 0 Reorder the terms: -30 + 5x2 = 0 Solving -30 + 5x2 = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '30' to each side of the equation. -30 + 30 + 5x2 = 0 + 30 Combine like terms: -30 + 30 = 0 0 + 5x2 = 0 + 30 5x2 = 0 + 30 Combine like terms: 0 + 30 = 30 5x2 = 30 Divide each side by '5'. x2 = 6 Simplifying x2 = 6 Take the square root of each side: x = {-2.449489743, 2.449489743}

OpenStudy (anonymous):

So it's B? :)

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

Awesome! Thanks so much! Could you help me with a couple more please? :)

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

Solve the equation by completing the square. Round to the nearest tenth. x^2+8x=10 A. 1.1, 9.1 B. 1.1, -9.1 C. -1.1, 9.1 -1.1, -9.1

OpenStudy (anonymous):

I'm sorry... I can't help with that one. I keep getting x1 equals the square root of 26-4 and x2 equals negative the square root of 26-4.

OpenStudy (anonymous):

Hold on, let me check if I typed it in wrong :)

OpenStudy (anonymous):

Alright, thanks for helping! :)

OpenStudy (anonymous):

Your welcome

OpenStudy (anonymous):

Sorry I couldn't help with the last Question...

OpenStudy (anonymous):

I'm gonna ask another question to see if anyone can find the answer haha. It's alright! :) You did a lot for me :)